If a plane travels 170 miles per hour how much would it travel in 45 mins?

Answer by jim_thompson5910(35256)

45 min = 45/60 = 3/4 = 0.75 hr

d = r*t
d = 170*0.75
d = 127.5

The plane travels 127.5 miles